OpenStudy (anonymous):

If the path of a diver diving from a diving board is given by the model: h = - 0.44t2 + 2.2t + 10 where the h represents the height of the diver at t seconds after she leaves the board. Find the time (in seconds) that it takes for her to reach the maximum height for the dive.

OpenStudy (anonymous):

0.88 sec

OpenStudy (kropot72):

Differentiate h with respect to t as a first step. Can you do this?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Good. What do you get for dh/dt?

OpenStudy (anonymous):

nevermind i have no clue

OpenStudy (kropot72):

\[\frac{dh}{dt}=-0.88t+2.2\] If the derivative is put equal to zero we can solve to find t for the maximum height. -0.88t + 2.2 = 0 0.88t = 2.2 t = 2.2/0.88 = 2.5 seconds

OpenStudy (anonymous):

differentiate???

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

heck no max is at \(-\frac{b}{2a}=-\frac{2.2}{2\times -.44}=2.5\)

OpenStudy (anonymous):

i am going to bet ( i could certainly be mistaken) that @tbrooks3 is not taking a calc class for this. maybe i am wrong

OpenStudy (anonymous):

this is a algebra 1 class at my school

OpenStudy (anonymous):

that is what i thought. so "derivative" makes no sense to you right?

OpenStudy (anonymous):

right.. ive been told that soo many times today but had no clue

OpenStudy (kropot72):

Why did @tbrooks3 answer yes to "Can you do this?" Never mind.

OpenStudy (anonymous):

i had read something similar thru my class and got confused @kropot72

OpenStudy (anonymous):

that is what i thought. if you have something that looks like \[y=ax^2+bx+c\] or \[h=at^2+bt+c\] then max or min occurs at vertex, which you find by computing \(-\frac{b}{2a}\)

OpenStudy (kropot72):

@satellite73 Point taken. Thank you for guidance :)